7. What is the gravitational field strength at a distance of 60.0 km above the surface of the Earth?


1
Expert's answer
2021-07-28T09:09:03-0400

Gravitational field strength (in other words, gravitational acceleration) is given as follows:


"g = \dfrac{GM}{R^2}"

where "G = 6.674\times 10^{-11}\dfrac{m^3}{kg\cdot s^2}" is the gravitational constant, "M = 5.972\times 10^{24}kg" is the mass of the Earth, and "R = 6.371\times 10^{6}m + 0.06\times 10^{6}m = 6.431\times 10^{6}m" is the distance from the center of the Earth to the required point above the surface (radius plus 60 km).

Thus, obtain:


"g = \dfrac{6.674\times 10^{-11}\cdot 5.972\times 10^{24}}{(6.431\times 10^{6})^2} \approx 9.64m/s^2"

Answer. "g = 9.64m\/s^2".